Hydrogen peroxide decomposes to water and oxygen gas. The rate law is -Δ[H2O2]/Δt=k[H2O2]. If a sample with an initial concentration of 0.151 mol/L has a concentration of 0.128 mol/L after 153 minutes, what is the rate constant?
Answer #1

According to the rate, order is first order.

It's equation is:-

K = (2.303 / t) log (A°/At)

Where A° is initial concentration and At is concentration at time t.

K =( 2.303/153) log (0.151 / 0.128)

K = 1.08 × 10^-3 min^-1.